A cube of green jelly rests on a table. The jelly is pushed laterally with a
force of 30 N at it top surface and changes into a parallelogram shape
instead of sliding, with a lateral displacement of 0.125 m. Calculate the
shear strain induced in the cube of jelly.
1
Expert's answer
2020-03-16T13:17:50-0400

Here it is given that the lateral displacement of the tip of the vertical edge is 0.125 m. This corresponds to the edge turning through 30 degrees. So we have,


"\tan 30=\frac{0.125}{y}\to y=0.217\ m"

Thus the area of a face of the cube is,


"A=y^2=0.0468\ m^2"

Therefore the shearing stress is,


"\sigma=\frac{30}{0.0468}=641\frac{N}{m^2}"

The shearing strain is


"\theta=30\degree=\frac{\pi}{6}\ rad"